OpenStudy (anonymous):

Simplify 7 + 11 – 6 divided by 3 x 4 A.) 1 B.) 8 C.) 10 D.) 16

OpenStudy (anonymous):

12/12=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

that's your answer...do you know why?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

oh wait, i see where i made the mistake. Thanks lol

OpenStudy (anonymous):

1 is wrong O.o

OpenStudy (anonymous):

did it tell you the right answer?

OpenStudy (anonymous):

because the only answer that makes sense is 1 7+11-6=12 3 x 4=12 12/12=1